# Inexact Column Generation for Bayesian Network Structure Learning via Difference-of-Submodular Optimization

### v1

@@ 5.3 Result on Discrete Data
ALARM 0.08 0.06 0.10 0.16 0.09 0.07 0.09 0.11 83 82 77 67
Table 3 compares four methods on discrete datasets LUCAS ethz_lucas with $(n,N,d)=(12,2000,1)$, INSURANCE binder1997adaptive with $(n,N,d)=(27,20000,3.85)$ and ALARM beinlich1989alarm with $(n,N,d)=(37,20000,2.49)$. While GOBNILP demonstrates superior performance on small graphs, all the methods fail to produce structurally similar graphs for the larger ALARM dataset.
Although CG-DCA is an option for small discrete dataset, its performance degrades with larger networks. This is because the discrete BIC score contains a highly supermodular penalization term $(a_{i}-1)\prod_{j\in J}a_{j}$, making DCA less effective for DS optimization. Thus, development of an improved decomposition strategy for discrete scores (potentially other scores like BDeu buntine1991theory ) is essential to broaden the applicability of DCA to larger discrete Bayesian networks.
@@ 6 Conclusion
While CG-DCA is a viable approach for small graphs with discrete data, its scalability to larger graphs remains limited due to the high supermodularity induced by the penalty term in the discrete BIC function. To mitigate this limitation, one could explore alternative decomposition strategies for (potentially different) discrete scoring functions or develop new DS optimization techniques tailored to this specific computational challenge. Furthermore, building upon our efficient pricing heuristic, future research could be focused on developing exact pricing algorithms to enhance both the convergence guarantees and solution optimality of CG, by leveraging exact approaches for submodular lovasz1983submodular and supermodular nemhauser1981maximizing optimization.
@@ Appendix A Definitions
A.1 Scoring Functions
@@
For continuous data, we assume that variables $\{X_{i}\}_{i\in V}$ are mean-normalized such that $\mathbb{E}[X_{i}]=0$. We also assume that if $J$ is the parent set of node $i$, then $X_{i}\sim N(\alpha_{iJ}^{\top}\mathbf{X}_{J},\sigma_{i\leftarrow J}^{2})$ van20130 , where $\mathbf{X}_{J}=\{X_{j}\}_{j\in J}$. The parameters $\alpha_{iJ}$ and $\sigma_{i\leftarrow J}^{2}$ are unknown. The likelihood function defining $\texttt{score}_{i}(J)$ has included these parameters by their maximum likelihood estimator (MLE), thus its value only depends on $\text{pa}_{i}(G)$, not the other nodes. The resulting local score $\texttt{score}_{i}(J)$ in this context is
@@
$\hat{\sigma}^{2}_{i\leftarrow J}=\min_{\alpha\in\mathbb{R}^{|J|}}\mathbb{E}_{\hat{\mathbb{P}}}\left[(\alpha^{T}\mathbf{X}_{J}-X_{i})^{2}\right]$
denotes the empirical residual variance (under the empirical distribution $\hat{\mathbb{P}}$) of the linear regression predicting $X_{i}$ (with $\mathbb{E}[X_{i}]=0$) from the features $\mathbf{X}_{J}$. For simplicity, we will ignore the constant $-\frac{N}{2}(1+\log(2\pi))$ in our pricing problem optimization framework established based on Proposition 1.
@@
where $\mathcal{S}_{i}$ and $\mathcal{S}_{J}$ represent the sets of possible values that $X_{i}$ and $\mathbf{X}_{J}$ can take, respectively. The arity (i.e., the number of possible values it can take) of variable $X_{j}$ is denoted by $a_{j}$ for $j=1,\ldots,n$. The count function #$\#(x_{i},\mathbf{x}_{J})=\texttt{Count}(X_{i}=x_{i},\mathbf{X}_{J}=\mathbf{x}_{J})$ counts joint occurrences in the dataset, and #$\#(\mathbf{x}_{J})=\texttt{Count}(\mathbf{X}_{J}=\mathbf{x}_{J})$ provides the corresponding marginal counts.
@@ A.2 Submodular Function and Lovász Extension Function
We adopt standard definitions of the submodular set function and the Lovász extension lovasz1983submodular .
@@
A function is supermodular if $-f$ is submodular, and is modular if it is both submodular and supermodular.
@@
Define the nested subsets $S^{\sigma}_{k}=\{\sigma(1),\dots,\sigma(k)\}$ for $k\in\{1,\dots,d\}$, and $S^{\sigma}_{0}=\emptyset$. The Lovász extension $f^{L}:[0,1]^{n}\rightarrow\mathbb{R}$ of the set function $f$ at point $x$ is defined as
@@ Appendix B Row Generation and the Separation Problem
The exponential number of cluster constraints can be handled through row generation 10.5555/3020548.3020567 . This cutting plane approach starts with a restricted set $\hat{\mathcal{C}}$ and sequentially adds constraints violated by the current solution $x^{*}$. For an integer solution $x^{*}$, the separation problem identifies violated cluster constraints via cycle detection in the decoded graph from $x^{*}$ (using, e.g., depth-first search). For a fractional solution of RMLP, the separation problem identifies maximally violated cluster constraints through
@@ C.1 Proof of Proposition 1
$\hat{\sigma}^{2}_{i\leftarrow J}$ $=\min_{\alpha\in\mathbb{R}^{|J|}}\mathbb{E}_{\hat{\mathbb{P}}}[(\alpha^{T}\mathbf{X}_{J}-X_{i})^{2}]$
$=\min_{\alpha\in\mathbb{R}^{|J|}}\alpha^{T}\hat{\Sigma}_{J,J}\alpha-2\alpha^{T}\hat{\Sigma}_{J,i}+\hat{\Sigma}_{i,i}$
@@
where both $\log\text{det}(\hat{\Sigma}_{J,J})$ and $\log\text{det}(\hat{\Sigma}_{J\cup i,J\cup i})$ are submodular functions of set $J$ kelmans1983multiplicative . Function $\Lambda|J|+\lambda^{*}_{i}$ is modular in $J$.
@@
$z(J;\lambda^{*})=\underbrace{\log\text{det}(\hat{\Sigma}_{J\cup i,J\cup i})+\Lambda|J|+\sum_{\begin{subarray}{c}C\in\hat{\mathcal{C}}:\\ i\in C,~{}J\cap C\neq\emptyset\end{subarray}}\lambda_{C}^{*}+\lambda^{*}_{i}}_{\text{submodular}}-\underbrace{\log\text{det}(\hat{\Sigma}_{J,J})}_{\text{submodular}}.$
@@ C.2 Proof of Proposition 2
$\sum_{x_{i}\in S_{i}}\sum_{x_{J}\in S_{J}}\#(x_{i},\mathbf{x}_{J})\log\left(\frac{\#(x_{i},\mathbf{x}_{J})}{\#(\mathbf{x}_{J})}\right)$
$=$ $\sum_{x_{i}\in S_{i}}\sum_{x_{J}\in S_{J}}\#(x_{i},\mathbf{x}_{J})\log(\#(x_{i},\mathbf{x}_{J}))-\sum_{x_{J}\in S_{J}}\#(\mathbf{x}_{J})\log(\#(\mathbf{x}_{J}))$
$=$ $N\left[\sum_{x_{i}\in S_{i}}\sum_{x_{J}\in S_{J}}\hat{\mathbb{P}}(X_{i}=x_{i},X_{J}=x_{J})\log\hat{\mathbb{P}}(X_{i}=x_{i},X_{J}=x_{J})-\sum_{x_{J}\in S_{J}}\hat{\mathbb{P}}(X_{J}=x_{J})\log\hat{\mathbb{P}}(X_{J}=x_{J})\right]$
@@
$H(J):=-\sum_{x_{J}\in S_{J}}\hat{\mathbb{P}}(X_{J}=x_{J})\log\hat{\mathbb{P}}(X_{J}=x_{J})$
@@
$H(J\cup\{i\}):=-\sum_{x_{i}\in S_{i}}\sum_{x_{J}\in S_{J}}\hat{\mathbb{P}}(X_{i}=x_{i},X_{J}=x_{J})\log\hat{\mathbb{P}}(X_{i}=x_{i},X_{J}=x_{J})$
denote the entropy functions, which are known to be submodular fujishige1978polymatroidal .
@@
$z(J;\lambda^{*})=\underbrace{H(J\cup\{i\})+\sum_{\begin{subarray}{c}C:C\in\mathcal{C}\\ i\in C\\ C\cap J\neq\emptyset\end{subarray}}\lambda_{C}^{*}+\lambda^{*}_{i}}_{\text{submodular}}-\underbrace{\left(H(J)-\Lambda(a_{i}-1)\prod_{j\in J}a_{j}\right)}_{\text{submodular}}.$
@@ Appendix D Supplementary Numerical Results
Table 4: Performance of CG-DCA, GOBNILP (GI), HC, PC, and MMHC on Gaussian Datasets
@@
ALARM 0.08 0.06 0.13 0.10 0.16 0.09 0.07 0.15 0.09 0.11 83 82 81 77 67
Tables 4 and 5 present the comprehensive experimental results including the HC method, with $N\in\{5000,20000\}$ for Gaussian datasets.

### v2

@@ 5.3 Results on Discrete Data
ALARM 0.46 0.87 0.83 0.47 0.50 0.89 0.76 0.33 35 07 11 32
Table 3 compares four methods on discrete datasets LUCAS [23] with $(n,N,d)=(12,2000,1)$, INSURANCE [10] with $(n,N,d)=(27,20000,3.85)$ and ALARM [7] with $(n,N,d)=(37,20000,2.49)$. GOBNILP demonstrates superior performance in all three discrete datasets.
Although CG-DCA is an option for small discrete dataset, its performance degrades with larger networks. This is because the discrete BIC score contains a highly supermodular penalization term $(a_{i}-1)\prod_{j\in J}a_{j}$, making DCA less effective for DS optimization. Thus, development of an improved decomposition strategy for discrete scores (potentially other scores like BDeu [11]) is essential to broaden the applicability of DCA to larger discrete Bayesian networks.
@@ 6 Conclusion
While CG-DCA is a viable approach for small graphs with discrete data, its scalability to larger graphs remains limited due to the high supermodularity induced by the penalty term in the discrete BIC function. To mitigate this limitation, one could explore alternative decomposition strategies for (potentially different) discrete scoring functions or develop new DS optimization techniques tailored to this specific computational challenge. Furthermore, building upon our efficient pricing heuristic, future research could be focused on developing exact pricing algorithms to enhance both the convergence guarantees and solution optimality of CG, by leveraging exact approaches for submodular [36] and supermodular [39] optimization.
@@ Appendix A Definitions
A.1 The Data Matrix and Scoring Functions
We denote $D$ as the dataset, which is an $N\times n$ matrix where $N$ is the sample size and $n$ is the number of random variables (nodes in Bayesian Network). The entry $D_{ij}$ is the value of $j$-th variable $X_{j}$ in the $i$-th observation. For discrete labeled data, each label is mapped to a distinct numerical value.
@@ A.1 The Data Matrix and Scoring Functions
For continuous data, we assume that variables $\{X_{i}\}_{i\in V}$ are mean-normalized such that $\mathbb{E}[X_{i}]=0$. We also assume that if $J$ is the parent set of node $i$, then $X_{i}\sim N(\alpha_{iJ}^{\top}\mathbf{X}_{J},\sigma_{i\leftarrow J}^{2})$ [48], where $\mathbf{X}_{J}=\{X_{j}\}_{j\in J}$. The parameters $\alpha_{iJ}$ and $\sigma_{i\leftarrow J}^{2}$ are unknown. The likelihood function defining $\texttt{score}_{i}(J)$ has included these parameters by their maximum likelihood estimator, thus its value only depends on $J=\text{pa}_{i}(G)$. The resulting local score $\texttt{score}_{i}(J)$ in this context is
@@
$\hat{\sigma}^{2}_{i\leftarrow J}=\min_{\alpha\in\mathbb{R}^{|J|}}\mathbb{E}_{\hat{\mathbb{P}}}\left[(\alpha^{\top}\mathbf{X}_{J}-X_{i})^{2}\right]$
denotes the empirical residual variance (under the empirical distribution $\hat{\mathbb{P}}$) of the linear regression predicting $X_{i}$ (with $\mathbb{E}[X_{i}]=0$) from the features $\mathbf{X}_{J}$. For simplicity, we will ignore the constant $\frac{N}{2}(1+\log(2\pi))$ in our pricing problem optimization framework established based on Proposition 1.
@@
where $\mathcal{S}_{i}$ and $\mathcal{S}_{J}$ represent the sets of possible values that $X_{i}$ and $\mathbf{X}_{J}$ can take, respectively. The arity (i.e., the number of possible values it can take) of variable $X_{j}$ is denoted by $a_{j}$ for $j=1,\ldots,n$. The count function $(x_{i},\mathbf{x}_{J})=\texttt{Count}(X_{i}=x_{i},\mathbf{X}_{J}=\mathbf{x}_{J})$ counts joint occurrences in the dataset $D$, and $(\mathbf{x}_{J})=\texttt{Count}(\mathbf{X}_{J}=\mathbf{x}_{J})$ provides the corresponding marginal counts.
@@ A.2 Submodular Function and Lovász Extension Function
We adopt standard definitions of the submodular set function and the Lovász extension [36].
@@
A function is supermodular if f-f is submodular, and is modular if it is both submodular and supermodular.
@@
Define the nested subsets $S^{\sigma}_{k}=\{\sigma(1),\dots,\sigma(k)\}$ for $k\in\{1,\dots,d\}$, and $S^{\sigma}_{0}=\emptyset$. The Lovász extension $f^{L}:[0,1]^{d}\rightarrow\mathbb{R}$ of the set function $f$ at point $x$ is defined as
@@ Appendix B Row Generation and the Separation Problem
The exponential number of cluster constraints can be handled through row generation [15]. This cutting plane approach starts with a restricted set $\hat{\mathcal{C}}$ and sequentially adds constraints violated by the current solution $x^{*}$. For an integer solution $x^{*}$, the separation problem identifies violated cluster constraints via cycle detection in the decoded graph from $x^{*}$ (using, e.g., depth-first search). For a fractional solution of RMLP, the separation problem identifies maximally violated cluster constraints through
@@ C.1 Proof of Proposition 1
$\hat{\sigma}^{2}_{i\leftarrow J}$ $=\min_{\alpha\in\mathbb{R}^{|J|}}\mathbb{E}_{\hat{\mathbb{P}}}[(\alpha^{\top}\mathbf{X}_{J}-X_{i})^{2}]$
$=\min_{\alpha\in\mathbb{R}^{|J|}}\alpha^{\top}\hat{\Sigma}_{J,J}\alpha-2\alpha^{\top}\hat{\Sigma}_{J,i}+\hat{\Sigma}_{i,i}$
@@
where both $\log\text{det}(\hat{\Sigma}_{J,J})$ and $\log\text{det}(\hat{\Sigma}_{J\cup i,J\cup i})$ are submodular functions of set $J$ [34]. Function $\Lambda|J|+\lambda^{*}_{i}$ is modular in $J$.
@@
$z(J;\lambda^{*})=\underbrace{\frac{N}{2}\log\text{det}(\hat{\Sigma}_{J\cup i,J\cup i})+\Lambda|J|+\sum_{\begin{subarray}{c}C\in\hat{\mathcal{C}}:\\ i\in C,~J\cap C\neq\emptyset\end{subarray}}\lambda_{C}^{*}+\lambda^{*}_{i}}_{\text{submodular}}-\underbrace{\frac{N}{2}\log\text{det}(\hat{\Sigma}_{J,J})}_{\text{submodular}}.$
@@ C.2 Proof of Proposition 2
$\sum_{x_{i}\in S_{i}}\sum_{\mathbf{x}_{J}\in S_{J}}\#(x_{i},\mathbf{x}_{J})\log\left(\frac{\#(x_{i},\mathbf{x}_{J})}{\#(\mathbf{x}_{J})}\right)$
$=$ $\sum_{x_{i}\in S_{i}}\sum_{\mathbf{x}_{J}\in S_{J}}\#(x_{i},\mathbf{x}_{J})\log(\#(x_{i},\mathbf{x}_{J}))-\sum_{\mathbf{x}_{J}\in S_{J}}\#(\mathbf{x}_{J})\log(\#(\mathbf{x}_{J}))$
$=$ $N\left[\sum_{x_{i}\in S_{i}}\sum_{\mathbf{x}_{J}\in S_{J}}\hat{\mathbb{P}}(X_{i}=x_{i},\mathbf{X}_{J}=\mathbf{x}_{J})\log\hat{\mathbb{P}}(X_{i}=x_{i},\mathbf{X}_{J}=\mathbf{x}_{J})-\sum_{\mathbf{x}_{J}\in S_{J}}\hat{\mathbb{P}}(\mathbf{X}_{J}=\mathbf{x}_{J})\log\hat{\mathbb{P}}(\mathbf{X}_{J}=\mathbf{x}_{J})\right]$
@@
$H(J):=-\sum_{\mathbf{x}_{J}\in S_{J}}\hat{\mathbb{P}}(\mathbf{X}_{J}=\mathbf{x}_{J})\log\hat{\mathbb{P}}(\mathbf{X}_{J}=\mathbf{x}_{J})$
@@
$H(J\cup\{i\}):=-\sum_{x_{i}\in S_{i}}\sum_{\mathbf{x}_{J}\in S_{J}}\hat{\mathbb{P}}(X_{i}=x_{i},\mathbf{X}_{J}=\mathbf{x}_{J})\log\hat{\mathbb{P}}(X_{i}=x_{i},\mathbf{X}_{J}=\mathbf{x}_{J})$
denote the entropy functions, which are known to be submodular [24].
@@
$z(J;\lambda^{*})=\underbrace{N\cdot H(J\cup\{i\})+\sum_{\begin{subarray}{c}C:C\in\mathcal{C}\\ i\in C\\ C\cap J\neq\emptyset\end{subarray}}\lambda_{C}^{*}+\lambda^{*}_{i}}_{\text{submodular}}-\underbrace{\left(N\cdot H(J)-\Lambda(a_{i}-1)\prod_{j\in J}a_{j}\right)}_{\text{submodular}}.$
@@ Appendix D Supplementary Numerical Results
D.1 Comparison of DCA initialization methods
To evaluate the sensitivity of DCA initialization, we compare the BIC scores and time costs across three initialization methods (warm-start, random, and hybrid). The results are summarized in Table 4 and  5:
Table 4: BIC Score Comparison of the Three Initialization Approaches for DCA
$(n,N,d)$ hybrid random warmstart
(15, 05000,0.5) -101818.64 -101818.64 -101819.58
(15, 05000,1.0) -101954.63 -101939.75 -101975.70
(15, 05000,1.5) -101411.14 -101376.34 -101509.59
(15, 05000,2.0) -101646.26 -101556.54 -101608.25
(15, 20000,0.5) -406345.25 -406345.25 -406346.23
(15, 20000,1.0) -410415.68 -410415.68 -410382.11
(15, 20000,1.5) -405344.24 -405487.19 -405599.33
(15, 20000,2.0) -400909.42 -400861.51 -401710.97
(20, 05000,0.5) -136605.28 -136605.28 -136734.74
(20, 05000,1.0) -137222.13 -137212.97 -137363.27
(20, 05000,1.5) -136624.81 -136577.03 -136957.87
(20, 05000,2.0) -136849.69 -136610.04 -136750.44
(20, 20000,0.5) -535047.92 -535047.92 -535183.20
(20, 20000,1.0) -541850.16 -541887.02 -542168.27
(20, 20000,1.5) -538682.19 -538693.99 -539863.90
(20, 20000,2.0) -535301.38 -534115.87 -534939.03
(25, 05000,0.5) -170751.53 -170751.53 -170751.74
(25, 05000,1.0) -170279.79 -170296.43 -170302.01
(25, 05000,1.5) -171308.77 -171157.64 -171289.38
(25, 05000,2.0) -171333.89 -171105.48 -170760.08
(25, 20000,0.5) -677813.87 -677813.87 -677812.65
(25, 20000,1.0) -683475.44 -683189.30 -683826.26
(25, 20000,1.5) -676190.15 -676267.45 -676649.76
(25, 20000,2.0) -689041.33 -688690.44 -687147.17
*Each value is the average over ten independent instances.
Table 5: Runtime Comparison of the Three Initialization Approaches for DCA
$(n,N,d)$ hybrid (s) random (s) warmstart (s)
(15, 05000, 0.5) 0004.06 0004.04 001.17
(15, 05000, 1.0) 0025.76 0029.40 004.49
(15, 05000, 1.5) 0076.72 0087.43 010.00
(15, 05000, 2.0) 0170.39 0295.25 023.74
(15, 20000, 0.5) 0010.17 0010.94 003.51
(15, 20000, 1.0) 0036.71 0036.57 010.63
(15, 20000, 1.5) 0104.25 0109.73 023.95
(15, 20000, 2.0) 0229.80 0245.11 041.94
(20, 05000, 0.5) 0009.53 0009.60 002.42
(20, 05000, 1.0) 0084.34 0090.87 012.51
(20, 05000, 1.5) 0312.67 0430.37 048.78
(20, 05000, 2.0) 1039.53 1770.71 095.36
(20, 20000, 0.5) 0018.42 0018.54 008.71
(20, 20000, 1.0) 0133.69 0134.79 034.84
(20, 20000, 1.5) 0425.17 0504.29 444.37
(20, 20000, 2.0) 0997.99 1710.91 211.44
(25, 05000, 0.5) 0021.97 0022.35 003.87
(25, 05000, 1.0) 0261.84 0304.12 033.59
(25, 05000, 1.5) 0949.90 2434.03 131.69
(25, 05000, 2.0) 3054.63 6037.58 224.73
(25, 20000, 0.5) 0050.22 0053.13 012.48
(25, 20000, 1.0) 0404.43 0642.87 078.21
(25, 20000, 1.5) 1372.59 3688.33 227.15
(25, 20000, 2.0) 3189.98 6674.62 366.46
*Each value is the average over ten independent instances.
From the tables, we observe that
  • Random initialization yields highest (best among three) BIC scores in 15 out of 24 instances among three methods but requires the most time in 22 out of 24 instances.
  • Warm-start initialization achieves the lowest (worst among three) BIC scores in 16 out of 24 instances and is the most time-efficient in 23 out of 24 instances.
  • The hybrid approach strikes a balance between optimality and computational efficiency, since it focuses on local refinement around the current best pattern (through warm-start initialization) while building upon the foundation of global exploration (through random initialization) at the early stage.
D.2 Comparison of CG-DCA with Baselines
Table 6 presents the comprehensive experimental results including the HC method, with $N\in\{5000,20000\}$ for Gaussian datasets. For the three discrete datasets, Table  7 provides the score and time comparison for CG-DCA and GOBNILP, while Table 8 provides the graph comparisons of all baselines. The runtime performance of all baseline methods (implemented in R) on all datasets is summarized in Table 9 and  10.
Table 6: Performance of CG-DCA, GOBNILP (GI), HC, (Stable-)PC, and MMHC on Gaussian Datasets
@@ D.2 Comparison of CG-DCA with Baselines
Table 7: Score and Time Comparison of CG-DCA and GOBNILP on Discrete Datasets
Dataset BIC Score Gap (%) Time (seconds)
CG-DCA GOBNILP CG-DCA GOBNILP
LUCAS 0.03 0.00 0011.59 007.99
INSURANCE 0.75 0.00 1892.36 074.39
ALARM 2.85 0.00 4696.19 153.49
@@
ALARM 0.46 0.87 0.42 0.83 0.47 0.50 0.89 0.48 0.76 0.33 35 07 34 11 32
Table 9: Time Comparison of Stable PC, HC, and MMHC Algorithms on Gaussian Datasets
$(n,N,d)$ Stable PC (s) HC (s) MMHC (s)
(15, 05000, 0.5) 0.038 0.048 0.048
(15, 05000, 1.0) 0.035 0.054 0.047
(15, 05000, 1.5) 0.054 0.101 0.062
(15, 05000, 2.0) 0.095 0.098 0.092
(15, 20000, 0.5) 0.097 0.154 0.126
(15, 20000, 1.0) 0.121 0.196 0.152
(15, 20000, 1.5) 0.188 0.263 0.326
(15, 20000, 2.0) 0.306 0.420 0.889
(20, 05000, 0.5) 0.050 0.066 0.090
(20, 05000, 1.0) 0.105 0.159 0.112
(20, 05000, 1.5) 0.180 0.309 0.291
(20, 05000, 2.0) 0.266 0.342 0.425
(20, 20000, 0.5) 0.146 0.255 0.185
(20, 20000, 1.0) 0.279 0.481 0.438
(20, 20000, 1.5) 0.660 0.950 1.366
(20, 20000, 2.0) 1.327 2.331 3.634
(25, 05000, 0.5) 0.083 0.099 0.071
(25, 05000, 1.0) 0.095 0.133 0.098
(25, 05000, 1.5) 0.227 0.348 0.218
(25, 05000, 2.0) 0.273 0.403 0.351
(25, 20000, 0.5) 0.200 0.407 0.235
(25, 20000, 1.0) 0.288 0.637 0.583
(25, 20000, 1.5) 0.725 1.568 2.612
(25, 20000, 2.0) 1.158 2.087 2.940
*Each value is the average over ten independent instances.
Table 10: Time Comparison of Stable PC, HC, and MMHC Algorithms on Discrete Datasets
Dataset Stable PC (s) HC (s) MMHC (s)
LUCAS 0.08 0.08 0.08
INSURANCE 0.65 0.56 0.45
ALARM 0.61 0.75 0.53